Sheffield Company’s standard labor cost of producing one unit of Product DD is 3.9 hours at the rate of $13.2 per hour. During August, 40,800 hours of labor are incurred at a cost of $13.30 per hour to produce 10,300 units of Product DD. Compute the total labor variance. Total labor variance $ Compute the labor price and quantity variances. Labor price variance $ Labor quantity variance $ Compute the labor price and quantity variances, assuming the standard is 4.2 hours of direct labor at $13.50 per hour. Labor price variance $ Labor quantity variance $
Answer #1
a
Total labor variance 12396 Unfavorable =(40800*13.3)-(10300*3.9*13.2)
b
Labor price variance 4080 Unfavorable =40800*(13.3-13.2)
Labor quantity variance 8316 Unfavorable =13.2*(40800-10300*3.9)
c
Labor price variance 8160 Favorable =40800*(13.3-13.5)
Labor quantity variance 33210 Favorable =13.5*(40800-10300*4.2)